### v1

# A Study of Library Usage in Agent-Authored Pull Requests

Abstract.
Coding agents are becoming increasingly capable of completing end-to-end software engineering workflows that previously required a human developer, including raising pull requests (PRs) to propose their changes. However, we still know little about how these agents use libraries when generating code, a core part of real-world software development. To fill this gap, we study 26,760 agent-authored PRs from the AIDev dataset to examine three questions: how often do agents import libraries, how often do they introduce new dependencies (and with what versioning), and which specific libraries do they choose? We find that agents often import libraries (29.5% of PRs) but rarely add new dependencies (1.3% of PRs); and when they do, they follow strong versioning practices (75.0% specify a version), an improvement on direct LLM usage where versions are rarely mentioned. Generally, agents draw from a surprisingly diverse set of external libraries, contrasting with the limited “library preferences” seen in prior non-agentic LLM studies. Our results offer an early empirical view into how AI coding agents interact with today’s software ecosystems.
1. Introduction
Large language models (LLMs) have rapidly reshaped the way developers write code, moving beyond single-line completions to generating whole functions or creating new projects (chenEvaluatingLargeLanguage2021; jiangSurveyLargeLanguage2024). More recently, the field has seen a shift from human-centric code completion to agentic workflows in which LLMs plan, act, and call tools to solve more realistic software engineering tasks (liuLargeLanguageModelBased2025). These agentic systems—now widely available in production tools (shrivastavaComparativeFeaturesetAnalysis2025)—can complete end-to-end workflows that previously required a human developer. Crucially, these workflows include raising pull requests (PRs), allowing agents to propose substantial code changes rather than the narrow dependency updates handled by earlier bots (heAutomatingDependencyUpdates2023a).
For agents to contribute meaningfully to real projects, they must be able to work effectively with external libraries. This is central to modern software development (somervilleSoftwareEngineeringGlobal2016), where reusing established libraries improves efficiency and avoids reinventing well-tested functionality (xuWhyReinventingWheels2020). However, prior work shows that LLMs in non-agentic settings often struggle here: hallucinating package names (twistLibraryHallucinationsLLMs2025), calling deprecated APIs (wangLLMsMeetLibrary2025), or omitting version constraints altogether (rajRoleLibraryVersions2024a). LLMs often rely on retrieval-based solutions to overcome these issues (jainMitigatingCodeLLM2024; wangExploraCoderAdvancingCode2025), but agents have the autonomy to bypass this additional infrastructure. What remains unclear is whether agentic workflows will actually use external libraries—and if they do, whether this autonomy helps mitigate existing problems or simply reproduces them at scale.
To begin to fill this gap, we present the first focused empirical study of how AI coding agents use libraries in practice, by mining real agent-authored PRs from the AIDev dataset (liRiseAITeammates2025). Specifically, we analyse file diffs (both code and dependency manifests) from 26,760 PRs across four languages (TypeScript, Python, Go, and C#), to answer three research questions (RQs):
RQ1 Library Usage How often do agents import libraries when generating code, and do they favour built-in or external libraries?
RQ2 New Dependencies How often do agents introduce new dependencies into a repository, and do they specify versions?
RQ3 Choosing Libraries Which specific libraries do agents choose, and what functional roles do they play?
Our findings are concise but instructive. (1) Agents often import libraries (29.5% of PRs) but rarely add entirely new dependencies (1.3% of PRs). (2) Agentic workflows encourage far better version hygiene than traditional LLM prompting: agents specify versions for 75.0% of added dependencies, whereas LLMs would only mention versions in $\approx$9% library-related interactions (rajRoleLibraryVersions2024a). (3) Agents use a surprisingly diverse range of libraries, importing 3,988 distinct external packages across 7,888 PRs—a sharp contrast to the narrow “library preferences” previously observed in more self-contained LLM tasks (twist-arxiv-llm-code-bias).
Our contributions are as follows:
  1. (1)
    We conduct the first empirical study of agentic library usage by mining 26,760 agent-authored PRs across four languages.
  2. (2)
    We find that agents often import libraries but rarely introduce new dependencies, and when they do, their versioning practices are unexpectedly strong.
  3. (3)
    We release our dataset processing and analysis code to support reproducibility and future research in this area:
    https://github.com/itsluketwist/agent-library-usage/
Table 1. Language-specific details for data extraction. This table summarises the code files, dependency manifests, and representative import/dependency syntax used to identify library usage across all four languages.
Programming Language Code File Extensions Example Import Statements Dependency Manifests Example Dependencies TypeScript (typescriptDocumentationModules2025) .ts, .tsx, .js, .jsx, .mjs, .cjs import { useEffect } from ’react’; import { Class } from ’@angular/core’; package.json "lodash": $\sim$4.17.21", Python (python5ImportSystem2025) .py import numpy as np from django.db import models requirements.txt requests pandas<2.0 pyproject.toml "pytest" "numpy==2.79", Go (goPackagesGoProgramming2025) .go import "fmt" import ( mux "github.com/gorilla/mux" ) go.mod require github.com/pkg/errors v0.9.1 C# (microsoftUsingDirectiveImport2025) .cs using Newtonsoft.Json; using Serilog; {name}.csproj packages.config <PackageReference     Include="Serilog" Version="2.12.0" /> paket.dependencies nuget Newtonsoft.Json 13.0.1
2. Dataset
To conduct our study, we use the AIDev dataset (liRiseAITeammates2025), a large-scale corpus of almost a million pull requests (PRs) from over 116,000 real GitHub repositories, authored by five distinct LLM-based coding agents (Claude Code, Cursor, Devin, Copilot, OpenAI Codex). Specifically, we use the November 2025 snapshot (commit eee0408 (hao-liHaoliAIDevEee0408a277826d88fc0ca5fa07d2fc325c96af12025)).
For our analysis, we use the AIDev-Pop subset that is curated to repositories with over 100 GitHub stars and includes the full commit details; AIDev-Pop contains 33,596 PRs from 2,807 repositories. We use three tables: repository, containing metadata about each repository; pull_request, which provides high-level PR information; and pr_commit_details, which contains file-level diffs and commit content. The dataset provides diffs in the “combined diff format” (gitGitGitdiffDocumentation), allowing us to extract new content from each file by selecting lines that begin with whitespace followed by exactly one “+”.
We restrict our study to the four most common languages in the dataset—those with the largest number of repositories—to ensure both depth of analysis and sufficient statistical support. The chosen languages are TypeScript/JavaScript (combined due to nearly identical syntax and library ecosystems, and will be referred to as TypeScript throughout the paper), Python, Go and C#. We initially included Rust (fifth most popular language), but its PRs relied heavily on internal workspace crates that were indistinguishable from external imports, so we limited our study to the top four languages.
Ethical considerations
AIDev contains no personal or sensitive information, and all data is publicly accessible under the open-source license of its respective repository (typically MIT, Apache-2.0, or GPL variants). As our analysis only uses publicly available code, and does not involve any human-subject data, our use of the dataset does not raise privacy or ethical concerns.
Table 2. Coding agent library usage (RQ1, RQ2). Summary statistics from the analysis of 26,760 agent-authored pull requests, showing how often agents import libraries and introduce new project dependencies.
Metric TypeScript Python Go C# All languages General statistics Total repositories 840 530 242 220 1,832 Total pull requests 7,480 7,190 10,107 1,983 26,760 Total PRs merged 4,428 (59.2%) 5,468 (76.1%) 8,361 (82.7%) 1,146 (57.8%) 19,403 (72.5%) RQ1 Library Usage PRs importing any library 2,993 (40.0%) 2,720 (37.8%) 1,273 (12.6%) 902 (45.5%) 7,888 (29.5%) PRs importing standard library 828 (11.1%) 2,383 (33.1%) 1,244 (12.3%) 743 (37.5%) 5,198 (19.4%) PRs importing external libraries 2,847 (38.1%) 1,914 (26.6%) 324 (3.2%) 738 (37.2%) 5,823 (21.8%) Avg. standard library imports per PR 0.26 1.28 0.63 0.49 0.69 Avg. external library imports per PR 1.87 0.73 0.08 0.73 0.80 Unique external libraries imported 2,495 948 314 231 3,988 PRs with library imports merged 1,595 (53.3%) 1,768 (65.0%) 918 (72.1%) 467 (51.8%) 4,748 (60.2%) RQ2 New Dependencies PRs adding new dependencies 9 (0.1%) 200 (2.8%) 16 (0.2%) 120 (6.1%) 345 (1.3%) Avg. new dependencies per PR 0.01 0.17 0.02 0.30 0.08 Total new dependencies 95 1,237 203 593 2,128 Dependencies specifying version 95 (100.0%) 1,020 (82.5%) 203 (100.0%) 278 (46.9%) 1,596 (75.0%) Unique new dependencies 85 552 188 294 1,119 PRs with new dependencies merged 6 (66.7%) 131 (65.5%) 5 (31.2%) 56 (46.7%) 198 (57.4%)
3. Data Extraction Method
In this section, we outline how we extract the data used to answer each RQ. Full implementation details–including all parsing scripts and processing steps–are available in our public GitHub repository.
3.1. Method for RQ1 Library Usage
We begin by investigating how inclined agents are to make use of libraries in their code. We recognise library usage only when imports appear in agent-generated code, because imports that already existed in a file prior to the agent’s edit do not reflect the agent’s preferences; and with only partial file snapshots provided in diffs, downstream function calls are often ambiguous without the corresponding import statement. We use file extensions to identify code files for each programming language, as defined by GitHub Linguist (github-linguistLinguistLanguageSavant).
We implement language-specific regex parsers to automatically extract import statements from those files, taking care to distinguish relative imports (from within the project itself). All extraction is supported by unit tests to ensure reliability. Information on the file extensions we consider, along with example imports, is provided in Table 1. Each library is categorised as either standard (built-in, as defined in the language documentation) or external (installed via a third-party package manager). This allows us to measure how heavily agents rely on external ecosystems versus native language capabilities.
3.2. Method for RQ2 New Dependencies
Next, we examine how often agents introduce new libraries into repositories. Here, we are particularly interested in the quality of these additions—specifically, whether agents are explicit with version constraints—because prior work shows that LLMs in non-agentic workflows rarely discuss versions (rajRoleLibraryVersions2024a), despite the known risks. For each language, we identify its corresponding dependency manifest files (for example, “requirements.txt” for Python, “package.json” for TypeScript) following the GitHub Docs (githubDependencyGraphSupported); we ignore lock files, which contain large numbers of second-hand dependencies not explicitly used within the repository.
We then implement manifest-specific regex parsers to automatically identify new dependencies—including any explicit version specifiers (for example, “==1.2” or “¿=1.0”)—again with unit tests to ensure reliability. The manifest files we consider, and example dependency definitions, are listed in Table 1. This enables us to quantify how often agents introduce completely new libraries and how reliably they do so.
3.3. Method for RQ3 Choosing Libraries
Finally, we examine the specific libraries agents choose and the functional roles they play. For each language, we take the set of libraries identified in RQ1 Library Usage and RQ2 New Dependencies and compute simple frequency statistics to surface the most commonly used or newly introduced libraries. We then perform manual analysis on the top ten libraries for each language and RQ to determine the areas that agents are most likely to use external libraries.
4. Findings & Discussion
We analyse a total of 26,760 agent-authored pull requests across 1,832 repositories and the four most common languages in the AIDev dataset. Here, we present and discuss our findings.
4.1. Findings for RQ1 Library Usage
For this RQ, we extract all libraries imported by agents and examine how often they rely on built-in standard libraries versus external third-party dependencies. The full breakdown is shown in Table 2.
Agents often import libraries when generating code, with 29.5% of PRs importing at least one library, and an average of 1.5 libraries imported across all PRs. However, the extent varies substantially by language: agents lean most heavily on libraries when working in TypeScript and C# (40.0% and 45.5% of PRs), while Go is a clear outlier, importing libraries in just 12.6% of PRs, suggesting that agents mostly modify the Go code without using additional tooling. This shows up again when looking at average imports per PR: both TypeScript and Python exceed two libraries per PR, whereas Go averages below one. Although reusing libraries is a core skill for professional developers, there is no established baseline for how often this should occur, making it difficult to assess whether these rates indicate overuse, underuse, or neither.
Agents rely strongly on the standard library in Python and C# (33.1% and 37.5% of PRs), but their use of external libraries in those languages is comparatively well balanced (26.6% and 37.2%). In TypeScript, the pattern is very different: agents overwhelmingly favour external packages (38.1% of PRs) over the standard library (11.1%), suggesting a strong inclination toward high-level tooling. Once again Go is the exception, using the standard library almost four times more often than external dependencies.
Finally, external library diversity is also high: agents collectively import 3,988 unique libraries. This contrasts sharply with earlier work showing that LLMs typically exhibit low library diversity (twist-arxiv-llm-code-bias). It suggests that giving agents rich project-level context nudges them toward much broader library use. While it may seem surprising that TypeScript alone contributes more than half of all unique libraries, this is entirely consistent with the ecosystem’s well-known fragmentation and reliance on micro-packages (kulaImpactMicroPackagesEmpirical2017).
RQ1 Summary
Agents often import libraries for TypeScript (40.0% of PRs), Python (37.8%) and C# (45.5%); and rarely for Go (12.6%). For TypeScript, agents rely on external libraries over built-in, for Python and C# we observe more balanced library usage, and for Go the standard library is highly preferred.
Table 3. Libraries used by agents (RQ3). Top external libraries imported and top newly added dependencies by language, with number of PRs. The lists reveal the functional roles (testing, scaffolding, UI, data, API clients) that agents most commonly serve.
Metric TypeScript Python Go C# Top 10 external library imports react (1013) vitest (483) zod (313) next (303) lucide-react (261) @testing-library/react (136) @tanstack/react-query (111) lodash (107) drizzle-orm (97) chai (92) pytest (812) pydantic (279) crewai (264) requests (125) numpy (125) fastapi (114) openai (65) pandas (60) httpx (53) yaml (52) github.com/stretchr/testify (108) github.com/smacker/go-tree-sitter (21) github.com/spf13/cobra (20) github.com/cometbft/cometbft (13) github.com/microsoft/typescript-go (13) github.com/tliron/glsp (13) go.uber.org/zap (12) github.com/gofiber/fiber (11) k8s.io/apimachinery/pkg (10) github.com/onsi/gomega (9) Avalonia (275) Xunit (243) Dock (131) Aspire (55) Azure (50) NUnit (48) FluentAssertions (30) ReactiveUI (28) Newtonsoft (27) Svg (27) Top 10 new dependencies typescript (3) vite (3) husky (2) @biomejs/biome (2) zod (2) prettier (2) @types/node (2) three (2) @electron-forge/plugin-auto-unpack-natives (1) eslint-plugin-react-compiler (1) requests (24) openai-agents (22) openai (21) numpy (21) pandas (18) pytest (14) python-dotenv (13) uvicorn (13) pydantic (12) aiohttp (12) github.com/golang/snappy (6) github.com/stretchr/testify (3) gopkg.in/yaml.v3 (3) github.com/klauspost/reedsolomon (3) github.com/gogo/protobuf (3) go.uber.org/atomic (2) github.com/smacker/go-tree-sitter (2) github.com/leonklingele/grouper (1) github.com/sashamelentyev/usestdlibvars (1) github.com/breml/bidichk (1) Microsoft.NET.Test.Sdk (26) xunit.runner.visualstudio (22) xunit (20) Moq (20) coverlet.collector (15) Microsoft.AspNetCore.Mvc.Testing (11) NUnit (9) NUnit3TestAdapter (9) Azure.Identity (7) Microsoft.Extensions.DependencyInjection (7)
4.2. Findings for RQ2 New Dependencies
For this RQ, we analyse dependency manifest files to investigate how often agents add new dependencies to projects. Full analysis is shown in Table 2.
Across all languages, dependency addition is extremely rare: only 1.3% of PRs introduce even a single new library. C# and Python stand out as the only languages where dependency addition is non-negligible (6.1% and 2.8% of PRs). This conservative behaviour contrasts with typical human workflows in large ecosystems—such as TypeScript and Python—where adding a dependency to reuse existing functionality is common practice (abdalkareemImpactUsingTrivial2020). Recent work shows that agents often struggle with installing packages and resolving environment conflicts (aroraSetupBenchAssessingSoftware2025), suggesting that their reluctance to add dependencies may reflect practical limitations rather than deliberate conservatism.
When agents do introduce new dependencies, they generally do so with good versioning practices. Across all languages, 75.0% of new dependencies include an explicit version constraint, with TypeScript and Go specifying versions in every instance. This contrasts with previous, non-agentic, LLM research that observes version constraints mentioned in only $\approx$9% of library-related discussions (rajRoleLibraryVersions2024a). Whilst not directly comparable, this does indicate that agentic setups may support more robust dependency management over direct LLM usage; likely because agents benefit from full project context and direct interaction with dependency manifests.
Finally, we observe substantial diversity in the dependencies added: 1,120 unique libraries across 2,129 total additions. This high ratio of unique packages suggests that agents do not default to a small, preferred set of libraries when adding new dependencies. Much of this diversity comes from a small number of diverse PRs that create entirely new projects and therefore introduce many dependencies at once; a pattern consistent with known strengths of modern LLMs in scaffolding initial project structures (rasnayakaEmpiricalStudyUsage2024).
RQ2 Summary
Agents rarely add new dependencies (1.3% of PRs), most commonly for C# (6.1%) and Python (2.8%). When they do add dependencies, they have strong version hygiene, specifying a constraint in 75.0% of cases.
4.3. Findings for RQ3 Choosing Libraries
For this RQ, we analyse trends in the libraries that agents most commonly import and add as new dependencies. Table 3 lists the most common libraries.
First, testing and development-tooling libraries appear across all languages. Although test frameworks are heavily-used, they make up only a minority of the distinct most-imported libraries (10–40%), with most of the imports instead being production-orientated. In contrast, newly added dependencies skew heavily toward testing and tooling, suggesting that when agents modify dependency manifests, they are mostly helping with project scaffolding rather than expanding runtime dependency graphs—an encouraging pattern from a security and maintenance perspective.
Second, language-specific choices follow ecosystem expectations. TypeScript imports focus on the front-end stack (react, next, lucide-react). Python shows a strong AI/data-science presence (crewai, openai, numpy, pandas) alongside ubiquitous HTTP clients (requests, httpx, aiohttp). Go leans toward infrastructure and cloud-native tooling (kubernetes packages, logging, CLI frameworks) alongside parser/protocol tools (tree-sitter, TypeScript compiler ports). C# imports include desktop/UI frameworks (Avalonia, Dock) and other Microsoft-ecosystem libraries, while new dependencies disproportionately configure test infrastructure (xUnit, NUnit, Moq). Overall, these trends show that agents generally choose sensible libraries for the given ecosystem.
RQ3 Summary
Across languages, agents import a broad set of production libraries but mostly add only test and developer tooling dependencies, indicating a conservative and tooling-oriented approach to dependency growth.
5. Threats to Validity
There are a few important limitations to consider when interpreting our findings. First, our results reflect behaviour in only a subset of AIDev, covering five deployed AI coding agents, four programming languages, and only high-quality repositories (100+ stars); other agents, languages, or repositories may exhibit different patterns. For internal validity, our extraction relies on regex parsing of file diffs, which can miss edge cases due to incomplete file context or complex syntax. Full abstract syntax tree parsing is not feasible because the diffs provide only partial files, so regex is the most practical option. We mitigated this threat by manually validating extraction from 100 diffs per language, and implementing comprehensive unit tests to ensure that expected syntax is correctly extracted.
6. Conclusion
We presented the first empirical study of how AI coding agents use libraries, analysing 26,760 agent-authored PRs across four languages. Our results show that agents frequently import libraries (29.5% of PRs) but rarely add new dependencies (1.3% of PRs), and when they do, versioning practices are strong. Agents also demonstrate surprisingly broad library diversity, with library imports centred around production code and new dependencies focussed on test code and developer tooling.
Looking ahead, the next step is to probe not just which libraries agents use, but how well they use them. For example, do they reproduce known LLM failure modes, such as hallucinating library names (twistLibraryHallucinationsLLMs2025) or relying on deprecated features (wangLLMsMeetLibrary2025)? A deeper analysis of the code they generate would also help establish whether agents sensibly reuse existing libraries or quietly reimplement functionality. As agentic workflows continue to mature, understanding these behaviours will be key to integrating them safely and effectively into real software development practices.